# Minute time scale variability in $\gamma$-ray flare of BL Lacertae

### v1

@@ 2.1 Fermi-LAT
Figure 1: Fermi-LAT light curves of BL Lacertae from MJD 60544 to MJD 60598. The first panel shows a 1-day binned light curve in the 0.1 - 500 GeV energy range. The second panel shows the temporal variation of the $\gamma$-ray photon index. The red line indicates the average PL index and the cyan line indicates the PL index mentioned in the 4FGL catalog. The third panel shows the 3-hr binned light curve. The blue lines indicate the Bayesian blocks and the red line indicates the average flux.
@@ 3 Results
Figure 2: Upper panel of each figure shows 5min-binned light curves of BL Lacertae in different orbits with TS value in the green bar and the lower panel shows the variation of PL index for that period. The red line is the peak fitted sum of the exponential function and their parameters are mentioned in Table 1.
Figure 3: Upper panel of each figure shows 2min-binned light curves of BL Lacertae in the same orbits mentioned in the 5-min binned light curve with TS value in the green bar and the lower panel shows the variation of PL index for that period.
@@ 3.1 $\gamma$-ray light curve
The spectral shape of BL Lacertae is defined as log-parabola in Fermi 4FGL catalog (Abdollahi et al., 2020) but we generated the light curve using FermiPy by modeling the spectra in each time bin with a power law (PL). The power law model is more suitable for this work as we are dealing with shortest scale $\gamma$-ray flux variation. The 1-day binned $\gamma$-ray light curve has been shown in the first panel of Figure 1 from MJD 60544 to MJD 60598. In this particular flaring period of BL Lacertae, we got the historical height daily binned flux (9.77 $\pm$ 0.37) $\times$ 10-6 ph cm-2 s-1 on 5th October 2024. The photon index calculated using PL for this brightest flux is 1.98 $\pm$ 0.03 which is slightly harder than the photon index value (2.143) mentioned in the Fermi 4FGL catalog.
Figure 4: 1-day binned $\gamma$-ray light curve in the different energy range of BL Lacertae. The energy range used to generate the light curve is mentioned in each panel.
We implemented the Bayesian blocks (BB) algorithm (Scargle et al., 2013) false-alarm probability (p0) of 0.05 to identify the flare in 3-hr binned $\gamma$-ray light curve with the help of a PYTHON package lightcurves [^5] https://github.com/swagner-astro/lightcurves (Wagner et al., 2022) shown in the last panel in Figure 1. With the help of the HOP algorithm, we identify the flare period with a condition ${}_{BB}\geq 3\bar{F}$, mean flux. Thus we have selected MJD 60578.98 to MJD 60598.00 as our main flaring period indicated in the black dotted line Figure 1.
We also generated 1-day binned light curve of BL Lacertae in 0.1-1 GeV, 1-50 GeV, and 50-500 GeV energy bands for that flaring period chosen previously with test statistic (TS) $>9$ following Pandey & Stalin 2022. As seen in Figure 4, light curves in 0.1-1 GeV and 1-50 GeV energy bands follow a similar trend of 0.1-500 GeV energy band whereas only one point is detected in 50-500 GeV energy band.
@@ 3.2 Minute-scale variability
We further investigate from MJD 60578.98142361 to MJD 60598 as we can see in the third panel of Figure 1, fluxes is around the 1 $\times$ 10-5 ph cm-2 s-1. We generated the light curves of this period with bin size equal to the orbital period ($\sim$ 95 min) of the Fermi-LAT shown in Figure 5. From the orbital binning $\gamma$-ray light curve, I have selected two sections S1 from MJD 60587.28003472 to MJD 60587.54623843 and S2 from MJD 60588.29855324 to MJD 60588.70364583 as they have higher TS value (TS $>250$) to generate 5 minutes binning $\gamma$-ray light curve. We have two chosen the orbit E1 lies within S1 and two orbits E2 and E3 from S2. We have estimated the shortest flux doubling/halving timescale from these light curves, as follows:
@@
where F($t_{1}$) and F($t_{2}$) are the flux values at times $t_{1}$ and $t_{2}$ , respectively, $\Delta t=t_{1}-t_{2}$ , and $t_{d}$ denotes the flux doubling/halving timescale. We observed the shortest flux halving timescale about 32.8 minutes from the orbital binned light curve. To study the temporal evaluation of flux, we fitted the peaks of a 5min-binned light curve with the sum of exponential in Figure 2 defined as
@@
where $T_{R}$ and $T_{D}$ are the rise and decay times of a particular flare and $F_{0}$ is the peak flux observed at time $t_{0}$. The shortest flux doubling/halving time, rise time, and decay time for each peak is defined in Table 1. We searched for minute-scale temporal variability of flux further by generating a 2 min-binned light curve of this period as we can see rapid flux variations with high TS values (e.g. Shukla et al. 2018; Pandey & Stalin 2022). The 2min-binned light curve is shown in Figure 3. We have also calculated the minimum flux doubling/halving timescale using equation 2 within the 2min-binned light curve. For E1, we found $t_{d}$approx 1.49 min where flux changed from (6.40 $\pm$ 1.48) $\times$ 10-6 to (1.62 $\pm$ 0.86) $\times$ 10-5 ph cm-2 s-1. For E2, we found $t_{d}$approx 1 min where flux changed from (2.46 $\pm$ 0.96) $\times$ 10-5 to (6.12 $\pm$ 0.49) $\times$ 10-6 ph cm-2 s-1.And for E3, we found $t_{d}$approx 1.48 min where flux changed from (7.40 $\pm$ 2.85) $\times$ 10-6 to (1.88 $\pm$ 0.89) $\times$ 10-5 ph cm-2 s-1.
Figure 5: Orbital binned $\gamma$-ray light curve of BL Lacertae with TS value in a light blue bar graph. S1 and S2 are the two selected periods for 5-min binning.
@@ 3.3 $\gamma$-ray spectral variability
Figure 7 shows the gamma-ray SED for each orbit. The log parabola model is used to fit the spectral points shown by the orange line in Figure 7 and the envelope shows the 1-$\sigma$ error. The $\gamma$-ray SED of E1 is modeled with index $\alpha$ = 1.93 $\pm$ 0.08 and a low curvature $\beta$approx 0.04. The $\gamma$-ray SED of E2 is modeled with index $\alpha$ = 1.93 $\pm$ 0.09 and a curvature $\beta$approx 0.16 and the SED of E3 is modeled with index $\alpha$ = 1.82 $\pm$ 0.07 and a low curvature $\beta$approx 0.04. This shows that orbit E1, E2, and E3 have spectra much closer to a power-law with a very small curvature. We also noticed that above 10 GeV mostly we have upper limits suggesting no high energy photons above 10 GeV are produced.
Figure 7: $\gamma$-ray SED for each orbit fitted with log-parabola model.
@@ 4 Discussion
Figure 8: The left figure is the X-ray light curve and the right figure shows the FLux vs PL index.
The central black hole mass of BL Lacertae is 3.8 $\times$ 108 M (Wu et al., 2009) and the corresponding event horizon light-cross time is $\sim$ 31 min. Our detected variability time-scale ($\sim$ 1 min) is much shorter than the event horizon light-cross time, suggesting that the enhancement in the $\gamma$-ray originated from a very compact region within the relativistic jet of BL Lacertae. This kind of ultra-fast variability can be explained by three scenarios: the origin of the flare may be a magnetospheric gap with a small volume near the event horizon of the black hole (Neronov & Aharonian, 2007), magnetic field reconnection in the highly magnetized jet when the emitter moves relativistically in the jet (Lyubarsky, 2005) and external objects (e.g. stars or clouds) may penetrate in the jet (Araudo et al., 2010). In this flaring case, the compact region near the innermost section of the jet is responsible for the GeV band flaring.

### v2

@@ 2.1 Fermi-LAT
Figure 1: (a) Fermi-LAT light curves of BL Lacertae from MJD 60544 to MJD 60598. The first panel shows a 1-day binned light curve in the 0.1 - 500 GeV energy range. The second panel shows the temporal variation of the $\gamma$-ray photon index. The red line indicates the average PL index and the cyan line indicates the PL index mentioned in the 4FGL catalog. The third panel shows the 3-hr binned light curve. The blue lines indicate the Bayesian blocks and the red line indicates the average flux. (b) 1-day binned $\gamma$-ray light curve in the different energy range of BL Lacertae. The energy range used to generate the light curve is mentioned in each panel.
@@ 3 Results
Figure 2: Upper panel of (a), (b) and (c) show 5min-binned light curves of BL Lacertae in different orbits with TS value in the green bar and the lower panel shows the variation of PL index for that period. The red line is the peak fitted sum of the exponential function and their parameters are mentioned in Table 1. Whereas (d), (e) and (f) show the same but with 2min-binned data.
@@ 3.1 $\gamma$-ray light curve
The spectral shape of BL Lacertae is defined as log-parabola in Fermi 4FGL catalog (Abdollahi et al., 2020) but we generated the light curve using FermiPy by modeling the spectra in each time bin with a power law (PL). The power law model is more suitable for this work as we are dealing with shortest scale $\gamma$-ray flux variation. The 1-day binned $\gamma$-ray light curve has been shown in the first panel of Figure 1(a) from MJD 60544 to MJD 60598. In this particular flaring period of BL Lacertae, we got the historical height daily binned flux (9.77 $\pm$ 0.37) $\times$ 10-6 ph cm-2 s-1 on 5th October 2024. The photon index calculated using PL for this brightest flux is 1.98 $\pm$ 0.03 which is slightly harder than the photon index value (2.143) mentioned in the Fermi 4FGL catalog.
We implemented the Bayesian blocks (BB) algorithm (Scargle et al., 2013) false-alarm probability (p0) of 0.05 to identify the flare in 3-hr binned $\gamma$-ray light curve with the help of a PYTHON package lightcurves [^5] https://github.com/swagner-astro/lightcurves (Wagner et al., 2022) shown in the last panel in Figure LABEL:fig:gamma_light_curve. With the help of the HOP algorithm, we identify the flare period with a condition ${}_{BB}\geq 3\bar{F}$, mean flux. Thus we have selected MJD 60578.98 to MJD 60598.00 as our main flaring period indicated in the black dotted line Figure 1(a).
We also generated 1-day binned light curve of BL Lacertae in 0.1-1 GeV, 1-50 GeV, and 50-500 GeV energy bands for that flaring period chosen previously with test statistic (TS) $>9$ following Pandey & Stalin 2022. As seen in Figure 1(b), light curves in 0.1-1 GeV and 1-50 GeV energy bands follow a similar trend of 0.1-500 GeV energy band whereas only one point is detected in 50-500 GeV energy band.
@@ 3.2 Minute-scale variability
We further investigate from MJD 60578.98 to MJD 60598.00 as we can see in the third panel of Figure 1(a), fluxes is around the 1 $\times$ 10-5 ph cm-2 s-1. We generated the light curves of this period with bin size equal to the orbital period ($\sim$ 95 min) using the Fermi-LAT data. From the orbital binning $\gamma$-ray light curve, we have selected two sections, S1 from MJD 60587.28003472 to MJD 60587.54623843 and S2 from MJD 60588.29855324 to MJD 60588.70364583 as they have higher TS value (TS $>250$) to generate 5 minutes binning $\gamma$-ray light curve. We have two chosen the orbit E1 lies within S1 and two orbits E2 and E3 from S2. We have estimated the shortest flux doubling/halving timescale from these light curves, as follows:
@@
where F($t_{1}$) and F($t_{2}$) are the flux values at times $t_{1}$ and $t_{2}$ , respectively, $\Delta t=t_{1}-t_{2}$ , and $t_{d}$ denotes the flux doubling/halving timescale. We observed the shortest flux halving timescale about 32.8 minutes from the orbital binned light curve. To study the temporal evaluation of flux, we fitted the peaks of a 5min-binned light curves with the sum of exponential in Figure 2 (a), (b) and (c) defined as
@@
where $T_{R}$ and $T_{D}$ are the rise and decay times of a particular flare and $F_{0}$ is the peak flux observed at time $t_{0}$. The shortest flux doubling/halving time, rise time, and decay time for each peak is defined in Table 1. We searched for minute-scale temporal variability of flux further by generating a 2 min-binned light curve of this period as we can see rapid flux variations with high TS values (e.g. Shukla et al. 2018; Pandey & Stalin 2022). The 2min-binned light curves are shown in Figure 2 (d), (e) and (f). We have also calculated the minimum flux doubling/halving timescale using equation 2 within the 2min-binned light curve. For E1, we found $t_{d}$approx 1.49 min where flux changed from (6.40 $\pm$ 1.48) $\times$ 10-6 to (1.62 $\pm$ 0.86) $\times$ 10-5 ph cm-2 s-1. For E2, we found $t_{d}$approx 1 min where flux changed from (2.46 $\pm$ 0.96) $\times$ 10-5 to (6.12 $\pm$ 0.49) $\times$ 10-6 ph cm-2 s-1.And for E3, we found $t_{d}$approx 1.48 min where flux changed from (7.40 $\pm$ 2.85) $\times$ 10-6 to (1.88 $\pm$ 0.89) $\times$ 10-5 ph cm-2 s-1.
@@ 3.3 $\gamma$-ray spectral variability
We have generated the gamma-ray SED for each orbit. The log parabola model is used to fit the spectral points in the SED. The $\gamma$-ray SED of E1 is modeled with index $\alpha$ = 1.93 $\pm$ 0.08 and a low curvature $\beta$approx 0.04. The $\gamma$-ray SED of E2 is modeled with index $\alpha$ = 1.93 $\pm$ 0.09 and a curvature $\beta$approx 0.16 and the SED of E3 is modeled with index $\alpha$ = 1.82 $\pm$ 0.07 and a low curvature $\beta$approx 0.04. This shows that orbit E1, E2, and E3 have spectra much closer to a power-law with a very small curvature. We also noticed that above 10 GeV mostly we have upper limits suggesting no high energy photons above 10 GeV are produced.
@@ 4 Discussion
Figure 4: (a) X-ray light curve (b) Flux vs PL index for X-ray observations.
The central black hole mass of BL Lacertae is 3.8 $\times$ 108 M (Wu et al., 2009) and the corresponding event horizon light-cross time is $\sim$ 31 min. Our detected variability time-scale ($\sim$ 1 min) is much shorter than the event horizon light-cross time, suggesting that the enhancement in the $\gamma$-ray originated from a very compact region within the relativistic jet of BL Lacertae. This kind of ultra-fast variability can be explained by three scenarios: the origin of the flare may be a magnetospheric gap with a small volume near the event horizon of the black hole (Neronov & Aharonian, 2007), magnetic field reconnection in the highly magnetized jet when the emitter moves relativistically in the jet (Lyubarsky, 2005) and external objects (e.g. stars or clouds) may penetrate in the jet (Araudo et al., 2010). In this flaring case, the compact region near the innermost section of the jet is responsible for the GeV band flaring. 2018ApJ...852..112K tried to explain the fast very high energy (VHE) variability of the Virgo Cluster radio galaxy M87 (NGC 4486) and the Perseus Cluster galaxy IC 310 (J0316+4119) with the gap-driven magnetospheric $\gamma$-ray emission from the rotating supermassive black holes. The day timescale variability of M87 has been explained by this magnetospheric gap near the event horizon whereas the 5-min timescale variability of IC 310 could not be explained by magnetospheric $\gamma$-ray emission. The minute-scale variability of BL Lacertae with black hole mass MBH $\approx$ 1.7 $\times$ 108 M (2014Natur.510..126Z) is less likely explained by this $\gamma$-ray emission from the magnetospheric gap near the event horizon.